OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 8 a.2 sqrt2 b.4 c.64

OpenStudy (whpalmer4):

Formula of a circle with the center at the origin is \(x^2+y^2=r^2\) where \(r\) is the radius. What value of \(r\) satisfies \(r^2 = 8\)?

OpenStudy (anonymous):

so \[x^2+y^2=8^2\]

OpenStudy (anonymous):

so its 64 right?

OpenStudy (whpalmer4):

does 64 * 64 = 8?

OpenStudy (whpalmer4):

The radius squared = 8. Alternatively, the radius = the square root of 8. You still like 64 as your answer?